# AI for Tabular Data

### v1

@@ Abstract
Recent foundation models for tabular data achieve strong task-specific performance via in-context learning. Nevertheless, they focus on direct prediction by encapsulating both representation learning and task-specific inference inside a single, resource-intensive network. This work specifically focuses on representation learning, i.e., on transferable, task-agnostic embeddings. We systematically evaluate task-agnostic representations from tabular foundation models (TabPFN and TabICL) alongside with classical feature engineering (TableVectorizer) across a variety of application tasks as outlier detection (ADBench) and supervised learning (TabArena Lite). We find that simple TableVectorizer features achieve comparable or superior performance while being up to three orders of magnitude faster than tabular foundation models. The code is available at https://github.com/ContactSoftwareAI/TabEmbedBench.
@@ 1 Introduction
Recent foundation models for tabular data, such as TabPFN hollmann2025tabpfn, TabICL tabicl, ConTextTab spinaci2025contexttab, and TabStar arazi2025tabstarf, leverage in-context learning to achieve strong task-specific performance on structured datasets. These models are trained on a huge number of synthetic datasets to capture underlying patterns in tabular data, enabling them to adapt quickly to new datasets through contextual information. However, unlike large language models (LLMs) that produce versatile, task-agnostic embeddings useful across diverse downstream applications - such as similarity search, retrieval, clustering, and few-shot transfer - these tabular models focus primarily on direct prediction rather than generating transferable representations.
Despite the significant progress in task-specific tabular learning, there remains a substantial gap in exploring task-agnostic embeddings for tabular data. From both research and industrial perspectives, such embeddings offer valuable flexibility by producing representations that can be reused across multiple tasks - including outlier detection, classification, and regression - without task-specific customization or retraining. Yet, to the best of the authors’ knowledge, there has not been a comprehensive evaluation comparing task-agnostic embedding models for tabular data.
This paper addresses this gap by systematically evaluating embeddings extracted from foundation models like TabPFN and TabICL as task-agnostic representations of individual data samples (table rows), enabling tasks like similarity search, outlier detection, regression and classification on a single table. They are a step towards applications such as cross-dataset search in data lakes, entity resolution, and comparing records across different sources. Therefore, this study focuses on row embeddings since individual records are the basic unit for most machine learning tasks, unlike column or table embeddings. Our primary contribution is to initiate discussion about the potential of general-purpose tabular embeddings, drawing parallels to the success of LLM embeddings. By highlighting both promise and current limitations we aim to establish task-agnostic tabular embeddings as a valuable research direction for the community.
@@ 2 Embedding Models
Learning effective task-agnostic representations of tabular data is challenging due to its heterogeneous nature, combining numerical and categorical features with complex dependencies that vary across datasets. While this initial study focuses on numerical and categorical tabular data, further research should incorporate other modalities such as text, see Section 5. We employ three approaches to obtain per-row tabular embeddings, spanning classical feature engineering to deep learned embeddings: TableVectorizer for statistical encoding, TabICL and TabPFN as tabular foundation models pre-trained on specific tasks yet capable of generating task-agnostic row embeddings.
@@
TableVectorizer from the skrub package skrub serves as a preprocessing baseline that converts tabular data into numerical representations through adaptive column-wise encoding, automatically selecting appropriate encoders (one-hot encoding, datetime decomposition, or dimensionality-reduced string encoding) based on each column’s characteristics before concatenating features into row embeddings. It is a non-learned, unsupervised method whose embedding dimension depends on the dataset characteristics (ranging from 4 to 166 in our benchmarks).
TabICL.
TabICL tabicl is a tabular foundation model that uses In-Context Learning for classification without parameter updates. The embeddings are extracted from the first stage, where a Set Transformer processes columns independently to create distribution-aware features. This is followed by a transformer with rotary positional embeddings that models inter-feature dependencies row-by-row. This process produces semantically rich representations via learnable [CLS] tokens. The embedding dimension is 512.
TabPFN.
TabPFN hollmann2025tabpfn, an additional foundation model that has been pre-trained on millions of synthetic datasets, employs a two-way attention mechanism that alternates between features and samples. Following the tabpfn-extensions approach tabpfn-extensions, we generate embeddings through feature masking. In this process, each feature is iteratively predicted from the remaining features using the TabPFNClassifier or TabPFNRegressor. We then extract the representations of each feature that are aggregated into final row embeddings by taking the average per component. The embedding dimension is 192.
@@ 3 Experiments
We evaluate embeddings on two benchmark suites: ADBench adbench for unsupervised outlier detection tasks and TabArena tabarena for supervised classification and regression tasks. From ADBench, we select only datasets labeled as tabular while excluding those from the image category, resulting in a curated collection of real-world tabular datasets. For TabArena, we use the TabArena Lite preset, which evaluates only the first fold of the first repeat for all tasks. This lightweight configuration reduces computational requirements while maintaining dataset diversity across 51 manually curated OpenML datasets. We impose upper bounds of 10,000 samples and 200 features per dataset to keep the computation time in a reasonable range. This results in 27 datasets each for ADBench and TabArena.
For each dataset, we generate embeddings and evaluate them using lightweight predictive algorithms (K-Nearest Neighbors, Multi-Layer Perceptrons with Optuna-based hyperparameter optimization optuna), and specialized outlier detection algorithms (Local Outlier Factor (LOF) localoutlier, Isolation Forest (IF) isolationforest, and DeepSVDD deepsvdd), details can be found in Appendix A. Following the respective benchmark protocols, TabArena Lite uses stratified train/test splits with single-fold evaluation, while ADBench outlier detection employs the full dataset approach.
As evaluation measures, we use the area under the receiver operating characteristic curve (AUC) for outlier detection and classification and the mean absolute percentage error (MAPE) for regression. The AUC score ranges from 0 to 1 and should be as large as possible, whereas the MAPE score is non-negative and should be as small as possible.
Table 1: Performance for outlier detection (AUC score $\pm$ standard deviation) for the three embedding models. The scores are averaged over 27 datasets.
Embedding Model Time (s) LOF IF DeepSVDD DeepSVDD-dyn
[…]
Table 2: Performance for binary classification (AUC score $\pm$ standard deviation) for the three embedding models. The scores are averaged over 19 datasets.
Embedding Model Time (s) KNNClassifier MLPClassifier
[…]
Table 3: Performance for multiclass classification (AUC score $\pm$ standard deviation) for the three embedding models. The scores are averaged over 6 datasets.
Embedding Model Time (s) KNNClassifier MLPClassifier
[…]
Table 4: Performance for regression (MAPE score $\pm$ standard deviation) for the three embedding models. The scores are averaged over 7 datasets.
Embedding Model Time (s) KNNRegressor MLPRegressor
[…]
4 Results and Discussion
The following discussion focuses on task-agnostic embeddings. Tabular foundation models such as TabPFN and TabICL are primarily constructed for prediction leveraging in-context learning. However, they also provide promising task-independent embeddings. Whenever we refer to TabPFN or TabICL, we mean the task-agnostic embeddings constructed by these models.
In the case of outlier detection, Table 1 shows that TableVectorizer embeddings outperform the ones of TabICL and TabPFN for every outlier algorithm. For LOF and IF, the task-agnostic embeddings of TabICL achieve better results than the ones of TabPFN, while the embeddings of the latter yield a higher AUC score for DeepSVDD and its dynamical variant. For the supervised tasks shown in Tables 23 and 4, TabICL performs best. Since the number of suited TabArena datasets was limited at the time of this investigation, general ranking conclusions for the three investigated embedding models cannot be drawn from these results. Nevertheless, for the used datasets it can be said, that TableVectorizer as simple feature engineering method keeps up or even outperforms tabular foundation models like TabICL and TabPFN when considering task-agnostic embeddings. Additionally, TableVectorizer constructs embeddings with the lowest dimension ($\leq 166$), while TabPFN requires 192 and TabICL even 512 as embedding dimension.
Beyond its performance, TableVectorizer offers a notable advantage in terms of embedding computation efficiency, requiring significantly less time than TabICL and TabPFN due to their more sophisticated model architecture. In particular, TabPFN requires a high amount of time since target-aware embeddings are computed for every column and are aggregated afterwards. In contrast, TabICL does not require customization as it naturally constructs task-agnostic embeddings, which are further processed by an in-context learning block when used directly for prediction.
A significant limitation of TableVectorizer is its variable embedding dimensionality, which is determined by the input dataset. This dataset-dependent dimensionality poses a challenge for applications requiring the joint embedding of diverse datasets into a unified representation space, which is a widespread scenario in industry. In contrast, models like TabPFN and TabICL employ a fixed, but significantly larger embedding size across all datasets.
5 Conclusion
This brief study demonstrates that simple feature engineering methods like TableVectorizer achieve competitive performance for task-agnostic embeddings while requiring substantially fewer computational resources in settings where sophisticated tabular foundation models cannot fully leverage their in-context learning capabilities. However, several open challenges remain for future work. First of all, our findings have to be confirmed on a wider range of datasets. Additionally, more embedding models should be included in this comparison. Although the embeddings are constructed in a task-independent way, to assess their task-agnostic quality, additional downstream tasks and metrics should evaluate the embeddings.
The ingestion of new rows raises fundamental questions about context selection for in-context learners: should entire datasets be re-embedded, or can a representative subset suffice? Distribution shift poses additional complications: If new data exhibit covariate shift or label drift, in-context models may produce embeddings misaligned with the original space.
Beyond these practical concerns, current methods primarily capture statistical patterns but lack semantic awareness. Recent works such as ConTextTab spinaci2025contexttab and TabSTAR arazi2025tabstarf demonstrate promising directions by incorporating textual metadata and target-aware representations, suggesting pathways towards semantically richer embeddings. Extending these approaches to relational databases represents a natural progression, where tabular foundation models could serve as building blocks for relational foundation models that learn representations across joined tables and foreign-key relationships.
Shared representation spaces across diverse datasets also warrant investigation: do models sufficiently differentiate between datasets when embedded jointly, or does cross-dataset contamination occur? Finally, assuming that semantically aware embeddings can be achieved, multimodal alignment becomes feasible, enabling tabular representations to be aligned with text, images, or other modalities for cross-modal retrieval and reasoning. Addressing these challenges will be essential for establishing task-agnostic tabular embeddings as a flexible, reusable primitive in the broader machine learning ecosystem.
Appendix A Experimental Details
All evaluators inherit from an abstract base class ensuring consistent training and prediction workflows. Below we describe the specific algorithms and their configurations.
A.1 Unsupervised Task: Outlier Detection
Local Outlier Factor (LOF).
@@ A.1 Unsupervised Task: Outlier Detection
A.2 Supervised Tasks: Classification and Regression
@@
We employ scikit-learn’s KNeighborsClassifier and KNeighborsRegressor with distance metrics euclidean and distance-based weighting scheme. For classification tasks, the model returns class probability predictions; for regression, it outputs continuous values. We iterate over the number of neighbors, i.e., n_neighbors $\in\{$5,10,15,20,25,30,35,40,45}\}, and record the optimal score per embedding model and dataset for evaluation.
@@
We implement a PyTorch-based MLP with Optuna hyperparameter optimization using Tree-structured Parzen Estimator (TPE) sampling. The search space encompasses: number of hidden layers (1-5 for classification, 1-3 for regression), hidden dimensions (32-512, log-scaled), dropout rates (0.0-0.5), learning rates (0.0001 to 0.01, log-scaled), batch sizes 16, 32, 64, 128, and training epochs (50-200). Models employ StandardScaler normalization, Adam optimization, and early stopping with 10-epoch patience. We conduct 50 optimization trials with 5-fold cross-validation.
Appendix B Detailed Results
In the following, we include tables and figures that show the results in more detail.
B.1 Outlier Detection
The AUC distribution for the three investigated embedding models and the different outlier detection algorithms are shown in Figure 1. The wide range of AUC values indicates that the difficulty of the investigated datasets strongly varies. The outlier ratio of the chosen 27 ADBench datasets ranges from 1% to 40%. Table 5 shows a breakdown of the AUC results with respect to the outlier ratio in the datasets. As can be seen, the performance deteriorates with increasing outlier ratio. The variance between the three investigated embedding models is most pronounced for small outlier ratios, it vanishes for large outlier ratios.
Figure 1: Distribution of the performance for outlier detection (AUC) for the three embedding models shown in Table 1.
Table 5: Performance for outlier detection (AUC score $\pm$ standard deviation) for the three embedding models, the chosen ADBench datasets are grouped by outlier ratio. The scores are averaged over the algorithms (DeepSVDD, DeepSVDD-dynamic, IsolationForest, LocalOutlierFactor) and datasets.
Outlier Ratio #Datasets TabICL TabPFN TableVectorizer

### v2

# Comparing Task-Agnostic Embedding Models for Tabular Data

@@ Abstract
Recent foundation models for tabular data achieve strong task-specific performance via in-context learning. Nevertheless, they focus on direct prediction by encapsulating both representation learning and task-specific inference inside a single, resource-intensive network. This work specifically focuses on representation learning, i.e., on transferable, task-agnostic embeddings. We systematically evaluate task-agnostic representations extracted from tabular foundation models (TabPFN, TabICL and TabSTAR) alongside classical feature engineering (TableVectorizer and a sphere model) across a variety of application tasks as outlier detection (ADBench) and supervised learning (TabArena Lite). We find that simple feature engineering methods achieve comparable or superior performance while requiring significantly less computational resources than tabular foundation models. The code is publicly available. [^1] https://anonymous.4open.science/r/ser34sgf4e5tabds23
@@ 1 Introduction
Recent foundation models for tabular data, such as TabPFN [7], TabICL [9] and TabSTAR [1], leverage in-context learning to achieve strong task-specific performance on structured datasets. These models are trained on a huge number of (synthetic) datasets to capture underlying patterns in tabular data, enabling them to adapt quickly to new datasets through contextual information. However, unlike large language models (LLMs) that produce versatile, task-agnostic embeddings useful across diverse downstream applications - such as similarity search, retrieval, clustering, and few-shot transfer - these tabular models focus primarily on direct prediction rather than generating transferable representations.
Despite the significant progress in task-specific tabular learning, there remains a substantial gap in exploring task-agnostic embeddings for tabular data. From both research and industrial perspectives, such embeddings offer valuable flexibility by producing representations that can be reused across multiple tasks - including outlier detection, classification and regression - without task-specific customization or retraining. Yet, to the best of the authors’ knowledge, there has not been a comprehensive evaluation comparing task-agnostic embedding models for tabular data.
This paper addresses this gap by systematically evaluating task-agnostic embeddings, enabling tasks like similarity search, outlier detection, regression and classification on a single table. They are a step towards applications such as cross-dataset search in data lakes, entity resolution, and comparing records across different sources. Therefore, this study focuses on row embeddings since individual records are the basic unit for most machine learning tasks, unlike column or table embeddings.
We consider two groups of embedding models: On the one hand embeddings extracted from foundation models like TabICL, TabPFN and TabSTAR as task-agnostic representations of individual data samples (table rows), and on the other hand classical feature engineering methods such as TableVectorizer. For the second group, we introduce a simple sphere model to overcome the limitation of a dataset-specific embedding dimension.
Our primary contribution is to initiate discussion about the potential of general-purpose tabular embeddings, drawing parallels to the success of LLM embeddings. By highlighting both promise and current limitations we aim to establish task-agnostic tabular embeddings as a valuable research direction for the community.
@@ 2 Embedding Models
Learning effective task-agnostic representations of tabular data is challenging due to its heterogeneous nature, combining numerical and categorical features with complex dependencies that vary across datasets. While this initial study focuses on numerical and categorical tabular data, further research should incorporate other modalities such as text, see Section 6. We only use models that can be applied to unseen datasets without further training. Thus, we employ five main approaches to obtain per-row tabular embeddings, spanning classical feature engineering to deep learned embeddings: TabICL, TabPFN and TabSTAR as tabular foundation models pre-trained on specific tasks yet capable of generating task-agnostic row embeddings, and TableVectorizer and a sphere model for statistical encoding.
TabICL.
TabICL [9] is a tabular foundation model that uses in-context learning for classification without parameter updates. The embeddings are extracted from the first stage, where a set transformer processes columns independently to create distribution-aware features. This is followed by a transformer with rotary positional embeddings that models inter-feature dependencies row-by-row. This process produces semantically rich representations via learnable [CLS] tokens. The embedding dimension is 512.
TabPFN.
TabPFNv2 [7], an additional foundation model that has been pre-trained on millions of synthetic datasets, employs a two-way attention mechanism that alternates between features and samples. We use three different methods to compute task-agnostic embeddings, each with dimension 192:
  • TabPFNColumns. Following the tabpfn-extensions approach [^2] https://github.com/PriorLabs/tabpfn-extensions, we generate embeddings through feature masking. In this process, each feature is iteratively predicted from the remaining features using the TabPFNClassifier or TabPFNRegressor depending on whether the feature is numerical or categorical. We then extract the representations of each feature that are aggregated into final row embeddings by taking the average per component.
  • TabPFNConstant. As a more resource-efficient way we artificially add a constant classification target, employ the TabPFNClassifier and extract the final embeddings.
  • TabPFNRandom. This approach is similar to the prior approach with an artificial target drawn from a standard normal distribution and using the TabPFNRegressor.
TabSTAR.
Although TabSTAR [1] is explicitly designed for target-aware representation, we can extract task-agnostic embeddings after the data is processed by the interaction encoder. With a single artificial regression target of 0 the statistical information of one row is represented in a 384-dimensional embedding. TabSTAR includes the semantics of the column names and is even capable of processing longer text with potential beyond categorical and numerical data.
@@
TableVectorizer from the skrub package [^3] https://github.com/skrub-data/skrub serves as a preprocessing baseline that converts tabular data into numerical representations through adaptive column-wise encoding, automatically selecting appropriate encoders (one-hot encoding, datetime decomposition, or dimensionality-reduced string encoding) based on each column’s characteristics before concatenating features into row embeddings. It is a non-learned, unsupervised method whose embedding dimension depends on the dataset characteristics (ranging from 4 to 166 in our benchmarks).
Sphere Model.
To overcome the limitation of a dataset-specific embedding dimension, we propose to embed all entities occuring in the table on the unit sphere in $d$ dimensions, where $d$ can be arbitrarily chosen. Assume that the table has $n$ rows and $m$ columns. In addition, without restriction we assume that the first $l$ columns are numerical and the last mlm-l columns are categorical. Ordinal categorical columns possess an intrinsic order and, hence, are treated as numerical. When we refer to categorical columns, we mean nominal categorical columns.
Let us first consider the numerical columns. As numerical columns possess an intrinsic order, we want to transfer this information to the embeddings. The idea is to map every numerical column $1\leq i\leq l$ to a half of a great circle on the unit sphere. For each numerical column, we independently sample a vector uniformly on the unit sphere, denoted by $u_{1},\ldots,u_{l}$. This vector serves as a starting point of the embedding great circle. Additionally we need a vector perpendicular to $u_{i}$ describing the direction of the great circle on the unit sphere. We choose this vector $\tilde{u}_{i}$ as $\tilde{u}_{i,2k}=u_{i,2k+1}$ and $\tilde{u}_{i,2k+1}=-u_{i,2k}$ for $1\leq k\leq\lfloor\frac{d}{2}\rfloor$. If $d$ is an odd number, we set the last component of $\tilde{u}_{i}$ to zero. We transform every column entry $e_{ij}$, $1\leq j\leq n$, to an angle in the range $0.1\cdot\pi,0.9\cdot\pi]$ by
$\alpha_{ij}=\left(0.1+0.8\cdot\frac{e_{ij}-\min_{1\leq j\leq n}e_{ij}}{\max_{1\leq j\leq n}e_{ij}-\min_{1\leq j\leq n}e_{ij}}\right)\cdot\pi.$
Finally, the embedding reads as
$v^{\text{num}}_{ij}=\sin(\alpha_{ij})\cdot u_{i}+\cos(\alpha_{ij})\cdot\tilde{u}_{i}.$
Hence, within each numerical column, the order of the corresponding embeddings aligns with the original order of the table entries, as $\alpha_{ij}$ preserves and transfers it.
As a (nominal) categorical column has no intrinsic order, all categories can be embedded independently. We do this by simply choosing a random uniformly distributed $d$-dimensional unit vector for each category. Note, that in high-dimensional spaces two randomly chosen vectors are nearly orthogonal; hence the probability of choosing close embeddings for different categories is nearly zero.
Finally, we construct row embeddings $h_{1},\ldots,h_{n}$ by calculating the sample mean of the entity embeddings.
Besides its simplicity, the embedding method offers a straightforward extension to unseen rows. For new categories, we simply draw an additional random category vector. Numerical values exceeding the original range still align with the original order. We initially do not use the full half circle $0,\pi]$ to cope with additional data outside the range. If the distance of new values to the original values is large, they could be scaled by a sigmoid function to the angle ranges $(0,0.1\cdot\pi]$ and $0.9\cdot\pi,\pi)$.
Furthermore, missing data is handled by mapping an empty table cell onto the zero vector, as it is equally far from any other entity embedding vector.
@@ 3 Experiments
We evaluate the quality of task-agnostic embeddings on two benchmark suites, use several task-specific predictive algorithms and evaluate the performance employing suited evaluation measures.
3.1 Datasets
@@
As benchmark datasets, we use ADBench [6] for unsupervised outlier detection tasks and TabArena [4] for supervised classification and regression tasks. From ADBench, we select only datasets labeled as tabular while excluding those from the image category, resulting in a curated collection of 47 real-world tabular datasets. For TabArena, we use the TabArena Lite preset, which evaluates only the first fold of the first repeat for all tasks. This lightweight configuration reduces computational requirements while maintaining dataset diversity across 51 manually curated OpenML datasets. We impose upper bounds of 15,000 samples and 500 features per dataset to keep the computation time in a reasonable range, resulting in 29 datasets for ADBench and 27 for TabArena.
3.2 Predictive algorithms
For ADBench datasets the task-agnostic embeddings are evaluated using specialized outlier detection algorithms, i.e. Local Outlier Factor [2], Isolation Forest [8], and DeepSVDD [10]:
Local Outlier Factor.
@@ 3.2 Predictive algorithms
The supervised tasks in the TabArena datasets are evaluated with K-Nearest Neighbors and Multi-Layer Perceptrons with Optuna [^4] https://optuna.org-based hyperparameter optimization (HPO):
@@
We employ scikit-learn’s KNeighborsClassifier / KNeighborsRegressor with distance metrics Euclidean and distance-based weighting scheme. For classification tasks, the model returns class probability predictions; for regression, it outputs continuous values. We iterate over the number of neighbors, i.e., n_neighbors $\in\{$5,10,15,20,25,30,35,40,45}\}, and record the optimal score per embedding model and dataset for evaluation.
@@
We implement a scikit-learn-based MLP with Optuna hyperparameter optimization using Tree-structured Parzen Estimator (TPE) sampling. The search space encompasses: number of hidden layers (1-5 for classification, 1-3 for regression), hidden dimensions (32-512, log-scaled), dropout rates (0.0-0.5), learning rates (0.0001 to 0.01, log-scaled) and batch sizes (16, 32, 64, 128). Models employ StandardScaler normalization, Adam optimization with 250 training epochs and early stopping with 10-epoch patience. We conduct 50 optimization trials with 5-fold cross-validation. For the optimal parameters, we use 1000 training epochs to obtain the final prediction.
All evaluators inherit from an abstract base class ensuring consistent training and prediction workflows. Following the respective benchmark protocols, TabArena Lite uses stratified train/test splits with single-fold evaluation, while ADBench outlier detection employs the full dataset approach.
3.3 Evaluation measures
We measure the prediction quality by the area under the receiver operating characteristic curve (AUC) for outlier detection and classification, with the one-vs-rest multiclass strategy. The AUC score ranges from 0 to 1 and should be as large as possible. In case of regression tasks, we use the mean absolute percentage error (MAPE). The MAPE score is non-negative and should be as small as possible.
To compare the performance of the embedding models across several predictive algorithms and datasets, we employ a critical difference diagram [3]. For each dataset and algorithm, the performance is ranked, where a lower rank indicates a better performance. The critical difference diagram plots the embedding models along a horizontal axis, ordered by their average rank. Black horizontal lines group algorithms with statistically non-significant differences in their ranks, as determined by a post-hoc test at significance level $\alpha$ = 0.05. Consequently, the diagram shows statistically equivalent performance groups and highlights significant differences between others.
In addition to the evaluation measures, we assess the performance of the embedding models by the required computation time (preprocessing and embedding computation, without evaluation). We conducted the experiments on a Google Cloud Platform virtual machine equipped with an NVIDIA A100 40GB GPU.
Figure 1: Critical difference diagram for outlier detection.
Figure 2: Distribution of the performance for Local Outlier Factor.
Figure 3: Distribution of the performance for Isolation Forest.
Figure 4: Distribution of the performance for DeepSVDD.
Figure 5: Distribution of the performance for DeepSVDD dynamic.
4 Results
The following discussion focuses on task-agnostic embeddings. Tabular foundation models such as TabICL, TabPFN and TabSTAR are primarily constructed for task-specific prediction leveraging in-context learning. However, they also provide promising task-independent embeddings. Whenever we refer to TabICL, TabPFN and TabSTAR, we mean the task-agnostic embeddings constructed by the specified model as described in Section 2. In contrast, we refer to the original end-to-end tabular foundation models as TabICLOriginal and TabPFNOriginal.
Fig. 1 shows the critical difference diagram summarizing the performance over all outlier tasks. From the statistical point of view, TableVectorizer and the sphere model with dimension 192 significantly outperform TabICL and the TabPFN versions. The sphere models, TableVectorizer and TabSTAR on the one hand, and the TabPFN versions alongside TabICL on the other hand are statistically indistinguishable.
To compare the performance for each of the four outlier detection algorithms separately, we plot the corresponding distribution of AUC scores via boxplots, see Figs. 2-5 (median: black line; mean: white square). The observations from Fig. 1 are roughly confirmed on the algorithmic level.
The supervised tasks, i.e. binary classification, multiclass classification and regression, are assessed by the critical difference diagram in Fig. 12. The embedding models cluster in two groups: One group containing the sphere models, TabSTAR, TabICL and TableVectorizer, the other one the TabPFN versions. The former group performs better than the latter, while the models within the first group are statistically indistinguishable. Considering only the rank order, TabSTAR and, more pronounced, TabICL perform better compared to outlier detection.
Figure 6: Distribution of the performance for binary classification evaluated via KNN-classifier.
Figure 7: Distribution of the performance for multiclass classification evaluated via KNN-classifier.
Figure 8: Distribution of the performance for regression via KNN. Note that for TableVectorizer and the TabPFN versions there are outliers that are not captured within the range of the vertical axis.
Figure 9: Distribution of the performance for binary classification evaluated via MLP-classifier.
Figure 10: Distribution of the performance for multiclass classification evaluated via MLP-classifier.
Figure 11: Distribution of the performance for regression evaluated via MLP-regressor. Note that for the TabPFN versions there are outliers that are not captured within the range of the vertical axis.